The difference of the squares of two consecutivee odd positive integers is 40. find the integers.
Answer by venugopalramana(3286)

any positive integer be x
even integer would be 2x
odd integer would be 2x+1
next odd integer would be 2x+1+2=2x+3
so difference in squares of these 2 numbers is 40
40= (2x+3)^2-(2x+1)^2=(2x+3+2x+1)(2x+3-2x-1)=(4x+4)2=8(x+1)
so x+1=40/8=5
x=5-1=4
so the 2 odd numbers are
2*4+1=9 and 9+2=11